Find an eqution of the line containg the given pair of points.
(-1,-2) and (-3,-7)
The eqution of the line in slope-intercept form is y=

Answer by mananth(16949)

x1 y1 x2 y2
-1 * -2 -3 -7

slope m = (y2-y1)/(x2-x1)
( -7 - -2 )/( -3 - -1 )
( -5 / -2 )
m= 5/ 2

Plug value of the slope and point ( -1,-2) in
Y = m x + b
-2.00 = -5/2 + b
b= -2 - 5/2
b= -9/ 2
So the equation will be
Y = 5/2 x -9/2
m.ananth@hotmail.ca